Let f be a degree two polynomial over R in n-variables and a = (a1, . . . , an) a
critical point of f. Show that there is a homogenous degree two polynomial Q
such that
f(x) − f(a) = Q(x − a) (x ∈ Rn
)
1
Expert's answer
2020-06-11T15:54:41-0400

Consider a polynomial of two variables "x_1, x_2"

"f(x_1,x_2)=c_{11}x_1^2+c_{22}x_2^2+c_{12}x_1x_2+\\
+c_1x_1+c_2x_2+c"

"a=(a_1,a_2)" - critical point of "f" :

"\frac{\partial f(a)}{\partial x_1}=2c_{11}a_1+c_{12}a_2+c_1=0\\
\frac{\partial f(a)}{\partial x_2}=2c_{22}a_2+c_{12}a_1+c_2=0"

Consider a polynomial

"f(x)-f(a)=c_{11}x_1^2+c_{22}x_2^2+c_{12}x_1x_2+\\
+c_1x_1+c_2x_2+c-c_{11}a_1^2-c_{22}a_2^2-c_{12}a_1a_2-\\
-c_1a_1-c_2a_2-c=\\
=c_{11}(x_1-a_1)(x_1+a_1)+\\
+c_{22}(x_2-a_2)(x_2+a_2)+\\
+c_{12}(x_1x_2-a_1x_2+a_1x_2-a_1a_2)+\\
+c_1(x_1-a_1)+c_2(x_2-a_2)=\\
=c_{11}(x_1-a_1)(x_1+a_1)+\\
+c_{22}(x_2-a_2)(x_2+a_2)+\\
+c_{12}x_2(x_1-a_1)+c_{12}a_1(x_2-a_2)+\\
+c_1(x_1-a_1)+c_2(x_2-a_2)=\\
=(x_1-a_1)\cdot\\
\cdot(c_{11}(x_1+a_1)+c_{12}x_2+c_1)+\\
+(x_2-a_2)\cdot\\
\cdot(c_{22}x_2(x_2+a_2)+c_{12}a_1+c_2)"

Consider

"c_{11}(x_1+a_1)+c_{12}x_2+c_1=\\
=c_{11}x_1+c_{12}x_2-c_{11}a_1-c_{12}a_2=\\
=c_{11}(x_1-a_1)+c_{12}(x_2-a_2)\\
c_{22}x_2(x_2+a_2)+c_{12}a_1+c_2=\\
=c_{22}x_2-c_{22}a_2=c_{22}(x_2-a_2)"

Then

"f(x)-f(a)=(x_1-a_1)\cdot\\
\cdot(c_{11}(x_1-a_1)+c_{12}(x_2-a_2))+\\
+(x_2-a_2)\cdot\\
\cdot c_{22}(x_2-a_2)=\\
=c_{11}(x_1-a_1)^2+\\
+c_{12}(x_1-a_1)(x_2-a_2)+\\
+ c_{22}(x_2-a_2)^2"

There is a homogenous degree two


Consider a polynomial of variables "x_1,x_2,...,x_n"

"f=\sum\limits_{i,j=1}^{n}c_{ij}x_ix_j+\sum\limits_{i=1}^{n}c_ix_i+c"

"a=(a_1,...,a_n)" - - critical point of "f" :

"\frac{\partial f(a)}{\partial x_1}=2c_{11}a_1+\\
+c_{12}a_2+...+c_{1n}a_n+c_1=0\\
...\\
\frac{\partial f(a)}{\partial x_n}=2c_{nn}a_n+\\
+c_{n1}a_1+...+c_{n,n-1}a_{n-1}+c_n=0"

"f(x)-f(a)=\\
=\sum\limits_{i,j=1}^{n}c_{ij}x_ix_j+\sum\limits_{i=1}^{n}c_ix_i+c-\\
-\sum\limits_{i,j=1}^{n}c_{ij}a_ia_j-\sum\limits_{i=1}^{n}c_ia_i-c=\\
=\sum\limits_{i,j=1}^{n}c_{ij}(x_i-a_i)(x_j-a_j)"

There is a homogenous degree two